# Introducing a Physics-informed Deep Learning Framework for Bridge Scour Prediction

### v2

@@ Abstract
This paper introduces scour physics-informed neural networks (SPINNs), a hybrid physics-data-driven framework for bridge scour prediction using deep learning. SPINNs are developed based on historical scour monitoring data and integrate physics-based empirical equations into neural networks as supplementary loss components. We incorporated three architectures: LSTM, CNN, and NLinear as the base data-driven model. Despite varying performance across different base models and bridges, SPINNs overall outperformed pure data-driven models. In some bridge cases, SPINN reduced forecasting errors by up to 50 percent. In this study, we also explored general models for bridge clusters, trained by aggregating datasets across multiple bridges in a region. The pure data-driven models mostly benefited from this approach, in particular bridges with limited data. However, bridge-specific SPINNs provided more accurate predictions than general SPINNs for almost all case studies. Also, the time-dependent empirical equations derived from SPINNs showed reasonable accuracy in estimating maximum scour depth, providing more accurate predictions compared to HEC-18. Comparing both SPINNs and pure deep learning models with traditional HEC-18 equation indicates substantial improvements in scour prediction accuracy. This study can pave the way for hybrid physics-machine learning methodologies to be implemented for bridge scour design and maintenance.
@@ 1   Introduction
Physics-based machine learning has accelerated as an emerging field in the past few years. The core idea is to incorporate physics law and constraints into the process of machine learning from data (Karpatne et al., 2017; Karniadakis et al., 2021). The hybrid physics-data-driven models have been introduced by a number of pioneering studies, referred to as physics-guided neural networks (PGNN) (Jia et al., 2021), physics-constrained neural networks (Raissi et al., 2019), and physics-informed neural networks (PINN) (Zhu et al., 2019). Although these terms have been alternately used in various recent studies (Faroughi et al., 2024). The incorporation of physical equations governing the response-input relationship, in the form of PDE/ODE or closed-form models within the loss function, is being increasingly referred to as physics-informed ML; hence, in this study, we stick to this term. PINNs have been particularly developed to solve PDEs in application to various problems in solid mechanics (Haghighat et al., 2021), porous media simulations (Chen et al., 2023; Amini et al., 2023), constitutive modelling, and soil consolidation (Eghbalian et al., 2023; Li et al., 2024a; Lan et al., 2024; Tian and Wang, 2023; Masi and Einav, 2024). Likewise PGNNs have found successful applications in structural analysis (Yu et al., 2020; Chen and Liu, 2021), geohazard assessment (Pei et al., 2023; Zhang et al., 2020) tunneling (Li et al., 2024b), fluid mechanics (Yousif et al., 2022), among other fields.
Building on previously developed algorithms (Yousefpour and Correa, 2023; Hashem and Yousefpour, 2024), this paper introduces hybrid physics-data-driven models for scour prediction. A novel physics-informed deep learning framework is developed by integrating empirical scour models, based on the commonly used HEC-18 form, with prominent deep learning algorithms. We introduce SPINNs, Scour Physics-Informed Neural Networks, which leverage the temporal feature extraction capability of recurrent neural networks, convolutional neural networks, and feed-forward neural networks, while respecting the physics-based empirical equations. The models are trained using historical scour monitoring data from a number of case-study bridges in Alaska. As a by-product of the physics-informed deep learning models, data-driven empirical models are introduced for scour prediction. A comparison between the pure data-driven models and the hybrid physics-data-driven models shows that the physics-informed DL models result in more accurate predictions and far better generalisation compared with the DL models with no physics-based constraints. Moreover, unlike data-driven DL models, which have been bridge/site-specific, the physics-informed DL models can provide a generalisable hybrid physics-data-driven model that can provide more reliable predictions for any bridges when trained with sufficient data.
@@ 2.2 Scour Data and Preprocessing
The scour monitoring data of four bridges in Alaska is used to train the scour prediction models in this study. This data includes time series of river bed, measured by sonar sensors at the bridge pier, and flow elevations, measured by stage sensors or gauges at the upstream, as well as river discharge, estimated based on measurement of flow velocity using velocimeter or Acoustic Doppler sensors. This data was provided by the Alaska Department of Transportation (DOT) and the US Geological Survey (USGS).
@@
2.3 Scour Physics-Informed Deep Learning Framework
Over the past few decades, various physics(hydraulics)-based empirical equations have been developed to estimate local scour depth based on physical factors such as flow depth, flow velocity, pier geometry, and riverbed material/geology, calibrated based on laboratory and field testing (Sheppard et al., 2011; Arneson et al., 2012; Kirby et al., 2015). Our hypothesis here is that integrating such semi-empirical physics-based equations into machine learning can lead to more competent hybrid physics-data-driven models. These models can learn from historical scour data while respecting the underlying physics, potentially achieving superior generalization performance compared to pure data-driven ML models. This new framework can also lead to the development of more reliable data-driven empirical equations through machine learning optimization process.
We introduce Scour Physics-informed Neural Network (SPINN) as a new physics-informed deep learning framework for scour prediction. This framework integrates an additional physical loss term into the objective function, guiding the deep learning models to learn from the specified empirical equation. The overall architecture of SPINN is depicted in Fig. 1.
@@ 2.3 Scour Physics-Informed Deep Learning Framework
The SPINN models are developed based on our pure data-driven LSTM and CNN scour forecast algorithms developed in Yousefpour and Correa (2023) and Hashem and Yousefpour (2024) for scour forecast based on historical sonar, stage, and discharge time series data. In addition, a simple DL linear architecture, NLinear is incorporated and compared with LSTM and CNN. The SPINN model processes time series data as input and output, including scour depth, $y_{s}$, flow depth, $y_{1}$, and discharge, $q$. The input sequence $X$, contains these three features over $M_{in}$ timesteps, while the output sequence, $Y_{s}$ contains only scour depth, $y_{s}$ over $M_{out}$ timesteps. The discrepancy (loss) between the target (actual) scour depth ($y_{i}$) and the predicted scour depth ($\hat{y}_{i}$) is quantified using the Mean Squared Error (MSE) over the length of the output sequence. As opposed to pure data-driven ML, the objective function is defined as the sum of two loss terms, a data-driven loss ($MSE_{D}$) and a physics-based loss ($MSE_{PHY}$), as described in Equations 1 and 2:
@@
Within this framework, $\hat{Y}_{s_{NN}}$ represents the output from the base NN model (pure ML), and $\hat{Y}_{s_{PHY}}$ denotes the predictions yielded by the empirical equation (calibrated through ML training). For this purpose, the $y_{1}$ and $q$ within the output timesteps ($M_{out}$) are used as input data for the physics empirical equation. By concurrently optimizing these two loss components, the SPINN model is supposed to achieve better predictive accuracy and generalizability.
In this study, we implement three SPINN variants, tailored by different forms of the empirical equations, ranging from widely accepted equations to a highly generalized form as described in the following sections. These equations involve latent parameters calibrated based on time series data throughout the training process. These parameters are optimized by Adam optimizer based on the total loss alongside the hidden parameters of the base NN model (Kingma and Ba, 2014).
@@ 2.3.1 SPINN with HEC18 Equation (SPINN-HEC18)
HEC-18, officially known as Hydraulic Engineering Circular No. 18, is a widely recognized engineering guideline published by the United States Federal Highway Administration (FHWA). This guideline provides recommendations on maximum scour depth prediction. Over time, HEC-18 has undergone iterative revisions, with the most current version documented in Arneson et al. (2012). The latest HEC-18 Equation (3) is shown in Equation (3):
@@
where $y_{s}$ is the equilibrium scour depth, $y_{1}$ is the flow depth, $a$ is the pier width, and $K_{1}$, $K_{2}$, $K_{3}$, are correction factors related to flow and pier as defined in Table 1. This equation is normally calibrated using datasets from field and laboratory measurements. These measurements represent different bridges with distinct scour characteristics. Unlike time series data, they are discretely collected, based on snapshots of scour conditions at different times and locations.
@@
In this paper, we use the time series data collected from continuous real-time sensor measurements on bridge piers, including riverbed elevation, flow elevation and discharge, $E_{bed}$, $E_{stage}$, and $q$. Therefore, $y_{s}$ and $y_{1}$ at each timestep can be derived by the following equations as shown in Fig. 2a:
@@
In addition, we include $p_{1}\in\left[-1,1\right]$ as another latent parameter adjusting the river and flow correction factor. Therefore, the empirical equation, ${\hat{Y}}_{s_{PHY}}$ can be written as:
@@ 2.3.2 SPINN with Time-Dependent HEC18 Equation (SPINN-TD)
Our dataset records show that the Alaskan bridges experience live-bed scour, where the riverbed goes through iterative cycles of scouring and filling (Ngo, 2018). For each occurrence of a scour event, the progression of live-bed scour can be outlined into two distinct phases, as illustrated in Fig. 3a. In the initial growth phase, local scour initiates from a depth of zero and gradually deepens. Upon reaching equilibrium depth, the local scour undergoes fluctuations, maintaining a relatively steady state. During this balanced phase, the flow dynamics lead to sediment deposition and erosion, resulting in a state of equilibrium. Furthermore, in the balanced phase, both upper and lower bounds of the scour depth become apparent. Consequently, various time-dependent equations (Sheppard et al., 2011; Liang et al., 2019) have been developed to approximate the upper bound of scour depth, as depicted in Fig. 3b.
@@
We define a time-dependent scour empirical model based on the maximum local scour depth equation of ($y_{s\_max}$) HEC18 and an exponential form to account for time-dependency as shown in Equation 8. During a scouring episode, the last timestep $M_{out}$ within the target sequence is considered close to the maximum scour depth in that episode, and the values of $y_{1_{max}}$ and $q_{max}$ are derived from this timestep.
@@
In this equation, ${\hat{y}}{st}$ is the predicted time-dependent local scour depth at timestep $t$, where $t\in\left\{0,1,2,\ldots,M_{out}-1\right\}$ represents the time steps of the output sequence. $T_{L}$ is the time lag between the maximum scour depth and the maximum flow, which is often observed in the scour process. This time lag is physically meaningful as it accounts for the delayed response of the scour depth to the flow conditions. $p_{1}$, $p_{2}$, and $p_{3}$ are latent parameters, where $p_{3}$ represents the decay rate in the exponential function, controlling how quickly the scour depth approaches the equilibrium state. These parameters are passed through a hyperbolic tangent (tanh) activation function, which constrains their values to the range $\left[-1,1\right]$. $T_{L}$ is set to be within the range [0, $\left[0,\ 2\times\left(M_{in}+M_{out}\right)\right]$. By setting the upper limit of $T_{L}$ to be twice the sum of $M_{in}$ and $M_{out}$, we ensure that the parameter has sufficient flexibility to learn the appropriate time lag between the maximum scour depth and the maximum flow, even if the lag extends beyond the output sequence length. To fit different scouring events throughout the time series, we consider the elevation in the first timestep of each training pair as $E_{ref}$, which is dynamically updated for different input and output data pairs.
@@
Given that the time-dependent equation mentioned above is specifically applicable to instances of scouring episodes, it is important to identify these episodes as illustrated in Fig. 4. The training workflow for a single data batch comprises several key steps. Initially, the bed elevation values of the first timestep within the mini-batch are picked as the reference level $E_{ref}$ for calculating flow depth and scour depth for each sequence, which are inputs for the NN model and the TD empirical equation. During this process, the sequences pass through the conditional term ($MSE_{PHY}$) to filter scour episodes, similar to SPINN-HEC18 models. The physics loss and Equation 8 are only applied to the scour episodes (i.e., positive scour depths).
@@
Similar to SPINN-HEC 18, the training is performed based on total loss for all mini-batches, including both scouring and filling episodes, to adjust the hidden parameters in the NN model and latent parameters in the TD empirical equation. The $MSE_{PHY}$ term is set to zero for filling episodes/sequences.
2.3.3 General Time-Dependent Equation (SPINN-GTD)
@@
SPINN-TD model has indeed incorporated the use of a time-dependent equation; however, it is important to note that using this equation requires an abundance of site-specific data, including pier geometry, pier shape, channel width, correction factors for bed condition, and flow attack angle. In order to develop a general model that does not require site-specific information, we introduce SPINN-GTD by eliminating all site-specific parameters from the empirical equation in the SPINN-TD (Equation 9), as outlined in Equation 9.
${\hat{y}}_{st}=p_{1}\left(y_{1_{max}}\right)^{\alpha}\left(\frac{q_{max}}{p_{2}}\right)^{\beta}\left(1-e^{-}\frac{t}{T_{L}}\right)$ (9)
@@
This approach enables the model to learn the features of a wider range of scour events in various bridges with different scour conditions, enhancing predictive accuracy and generalisation of the SPINN model to new bridge scour cases. Likewise, the generalised time-dependent equation calibrated through this SPINN model can potentially provide more accurate maximum scour depth predictions across a wide range of bridges and scour conditions.
2.3.4 Long Short-Term Memory Networks (LSTM)

### v3

# Physics-Inspired Deep Learning and Transferable Models for Bridge Scour Prediction

@@ Abstract
This paper introduces scour physics-inspired neural networks (SPINNs), a hybrid physics-data-driven framework for bridge scour prediction using deep learning. SPINNs integrate physics-based, empirical equations into deep neural networks and are trained using site-specific historical scour monitoring data. Long-short Term Memory Network (LSTM) and Convolutional Neural Network (CNN) are considered as the base deep learning (DL) models. We also explore transferable/general models, trained by aggregating datasets from a cluster of bridges, versus the site/bridge-specific models. Despite variation in performance, SPINNs outperformed pure data-driven models in the majority of cases. In some bridge cases, SPINN reduced forecasting errors by up to 50 percent. The pure data-driven models showed better transferability compared to hybrid models. The transferable DL models particularly proved effective for bridges with limited data. In addition, the calibrated time-dependent empirical equations derived from SPINNs showed great potential for maximum scour depth estimation, providing more accurate predictions compared to commonly used HEC-18 model. Comparing SPINNs with traditional empirical models indicates substantial improvements in scour prediction accuracy. This study can pave the way for further exploration of physics-inspired machine learning methods for scour prediction.
@@ 1   Introduction
Physics-based machine learning has accelerated as an emerging field in the past few years. The core idea is to incorporate physics law and constraints into the process of machine learning from data (Karpatne et al., 2017; Karniadakis et al., 2021). The hybrid physics-data-driven models have been introduced by a number of pioneering studies, referred to as physics-guided neural networks (PGNN) (Jia et al., 2021), physics-constrained neural networks (Raissi et al., 2019), and physics-informed neural networks (PINN) (Zhu et al., 2019). These terms have been used alternately in various recent studies (Faroughi et al., 2024). The incorporation of physical equations governing the input-response relationship, especially in the form of PDE/ODE within the loss function, is being increasingly referred to as physics-informed ML. PINNs have been particularly developed to solve PDEs in application to various problems in solid mechanics (Haghighat et al., 2021), porous media simulations (Chen et al., 2023; Amini et al., 2023), constitutive modelling, and soil consolidation (Eghbalian et al., 2023; Li et al., 2024a; Lan et al., 2024; Tian and Wang, 2023; Masi and Einav, 2024). Likewise PGNNs have found successful applications in structural analysis (Yu et al., 2020; Chen and Liu, 2021), geohazard assessment (Pei et al., 2023; Zhang et al., 2020) tunneling (Li et al., 2024b), fluid mechanics (Yousif et al., 2022), among other fields.
This paper introduces hybrid physics-data-driven models for scour prediction. A novel physics-inspired deep learning framework is developed by integrating semi-empirical scour models with prominent deep learning algorithms. As closed-form equations are used in this study, and to avoid misinterpretations, the term "physics-inspired neural networks" is adopted. We introduce SPINNs, Scour Physics-Inspired Neural Networks, which leverage the temporal feature extraction capability of recurrent neural networks and convolutional neural networks, while respecting the physics driving the scour process.
The SPINNs are trained using historical scour monitoring data (more than 15 years) from a number of case-study bridges in Alaska. As a by-product of SPINNs, new empirical equations are introduced that are calibrated through deep learning algorithms. These DL-calibrated empirical equations can have superior performance over locally calibrated empirical models (e.g., HEC-18).
We also explore the viability of transferable scour models; these general models are trained with collective monitoring data from a cluster of bridges within a region and implemented to predict scour for individual bridges. The transferability of the proposed data-driven and hybrid models are investigated across the case studies and their accuracy is compared to bridge/site-specific models.
@@ 2.2 Scour Data and Preprocessing
The scour monitoring data of four bridges in Alaska is used to train the scour prediction models in this study. This data includes time series of the river bed, measured by sonar sensors at the bridge pier, and flow elevations, measured by stage sensors or gauges at the upstream, as well as river discharge, estimated based on measurement of flow velocity using velocimeter or Acoustic Doppler sensors. This data was provided by the Alaska Department of Transportation (DOT) and the US Geological Survey (USGS).
@@
2.3 Physics-Inspired Deep Learning Framework
We introduce Scour Physics-inspired Neural Network (SPINN) as a new framework for scour prediction. This framework integrates an additional physical loss term using physics-based empirical equations. Over the past few decades, various semi-empirical equations have been developed to estimate local scour depth based on physical factors such as flow depth, flow velocity, pier geometry, and riverbed material/geology, calibrated based on laboratory and field testing (Sheppard et al., 2011; Arneson et al., 2012; Kirby et al., 2015). The overall architecture of SPINN is depicted in Fig. 1.
@@ 2.3 Physics-Inspired Deep Learning Framework
The SPINN models are developed based on our pure data-driven LSTM and CNN scour forecast algorithms developed in Yousefpour and Correa (2023) and Hashem and Yousefpour (2024) for scour forecast based on historical sonar, stage, and discharge time series data. The SPINN model receives the time series data in a number of small slices each including an input part and output part. The input sequence $X$, includes scour depth, $y_{s}$, flow depth, $y_{1}$, and discharge, $q$ over $M_{in}$ timesteps, while the output sequence, $Y_{s}$ contains only scour depth, $y_{s}$ over $M_{out}$ timesteps. The discrepancy (loss) between the target (actual) scour depth ($y_{i}$) and the predicted scour depth ($\hat{y}_{i}$) is quantified over the output sequence, using the Mean Squared Error (MSE). As opposed to pure data-driven ML, the objective function is defined as the sum of two loss terms, a data-driven loss ($MSE_{D}$) and a physics-based loss ($MSE_{PHY}$), as described in Equations 1 and 2:
@@
Within this framework, $\hat{Y}_{s_{NN}}$ represents the output from the base NN model (pure ML), and $\hat{Y}_{s_{PHY}}$ denotes the predictions yielded by the empirical equation (calibrated through ML training). For this purpose, the $y_{1}$ and $q$ within the output timesteps ($M_{out}$) are used as input data for the physics empirical equation. By concurrently optimizing these two loss components, the SPINN model learns both from the inherent relationship within data and also satisfies the underlying governing physical relationship dictated by the empirical equation.
In this study, we implement three SPINN variants, with three empirical equations, including the widely accepted HEC-18 equation to a modified time-dependent form as described in the following sections. These equations involve latent parameters calibrated based on time series data throughout the training process. These parameters are optimized by Adam optimizer based on the total loss alongside the hidden parameters of the base NN model (Kingma and Ba, 2014).
@@ 2.3.1 SPINN with HEC18 Equation (SPINN-HEC18)
HEC-18, officially known as Hydraulic Engineering Circular No. 18, is a widely recognized engineering guideline published by the United States Federal Highway Administration (FHWA). This guideline provides recommendations for the prediction of maximum depth of scour. Over time, HEC-18 has undergone iterative revisions, with the most current version documented in Arneson et al. (2012). The latest HEC-18 Equation (3) is shown in Equation (3):
@@
where $y_{s}$ is the equilibrium scour depth, $y_{1}$ is the flow depth, $a$ is the width of the pier, and $K_{1}$, $K_{2}$, $K_{3}$, are correction factors related to the conditions of the pier, the flow and the riverbed as defined in Table 1. This equation is predominantly calibrated using datasets from laboratory measurements.
@@
For training the SPINN model, we use time series of three features collected from continuous sensor measurements at bridge piers, including riverbed elevation, flow elevation and discharge, $E_{bed}$, $E_{stage}$, and $q$. Scour depth and flow depth, $y_{s}$ and $y_{1}$ at each timestep can be derived by the following equations, as shown in Fig. 2a:
@@
In addition, we include $p_{1}\in\left[-1,1\right]$ as another latent parameter that adjusts the combined correction factors (pier shape, flow direction, and riverbed). Therefore, the empirical equation, ${\hat{Y}}_{s_{PHY}}$ can be written as:
@@ 2.3.2 SPINN with Time-Dependent HEC18 Equation (SPINN-TD)
Our dataset records show that the Alaskan bridges experience live-bed scour, where the riverbed goes through iterative cycles of scouring and filling (Ngo, 2018). For each occurrence of a scour event, the progression of live-bed scour can be outlined into two distinct phases, as illustrated in Fig. 3a. In the initial growth phase, local scour begins from a depth of zero and gradually deepens. Upon reaching equilibrium depth, the local scour undergoes fluctuations, maintaining a relatively steady state. During this balanced phase, the flow dynamics lead to sediment deposition and erosion, resulting in a state of equilibrium. A number of time-dependent equations (Sheppard et al., 2011; Liang et al., 2019) have been developed to approximate the upper bound of scour depth, as depicted in Fig. 3b.
@@
We define a time-dependent empirical model of scour based on the maximum scour depth ($y_{s\_max}$) from HEC18 and an exponential decay function to account for time dependency as shown in Equation 8. In each data slice within a scour episode, the elevation at the first timestep is considered as $E_{ref}$, and the last timestep $M_{out}$ within the target sequence is considered as the maximum scour depth.
@@
In this equation, ${\hat{y}}{st}$ is the predicted time-dependent local scour depth at timestep $t$, where $t\in\left\{0,1,2,\ldots,M_{out}-1\right\}$ represents the time steps of the output sequence. $T_{L}$ is the time lag between the maximum scour depth and the maximum flow, which is often observed in the scour process. This time lag is physically meaningful as it accounts for the delayed response of the scour depth to the flow conditions. $p_{1}$, $p_{2}$, and $p_{3}$ are latent parameters, where $p_{3}$ represents the decay rate in the exponential function, controlling how quickly the scour depth approaches the equilibrium state. These parameters are passed through a hyperbolic tangent (tanh) activation function, which constrains their values to the range $\left[-1,1\right]$. $T_{L}$ is set to be within the range [0, $\left[0,\ 2\left(M_{in}+M_{out}\right)\right]$. By setting the upper limit of $T_{L}$ to be twice the sum of $M_{in}$ and $M_{out}$, we ensure that the parameter has sufficient flexibility to learn the appropriate time lag between the maximum scour depth and the maximum flow, even if the lag extends beyond the output sequence length.
@@
Given that the time-dependent equation mentioned above is specifically applicable to instances of scouring episodes, it is important to identify these episodes as illustrated in Fig. 4. The training workflow for a single data batch comprises several key steps. Initially, the bed elevation values of the first timestep within the mini-batch are picked as the reference level $E_{ref}$ for calculating flow depth and scour depth for each sequence, which are inputs for the NN model and the TD empirical equation. During this process, the sequences pass through the conditional term of physics loss ($MSE_{PHY}$) to filter scour episodes, similar to SPINN-HEC18 models. The physics loss is only applied to the scour episodes, therefore $MSE_{PHY}$ term is set to zero for filling episodes/sequences.
@@
Similar to SPINN-HEC18, the training is performed based on total loss for all mini-batches, including both scouring and filling episodes, to adjust the hidden parameters in the NN model and latent parameters in the TD empirical equation.
2.3.3 Transferable/General Time-Dependent Equation (SPINN-GTD)
@@
Both SPINN-HEC18 and SPINN-TD are site/bridge-specific models and require abundance of site-specific data, including pier geometry, and correction factors for bed condition, and flow attack angle. In order to develop a general model that is transferable across multiple bridges within a region (a cluster of bridges with similarities), we introduce SPINN-GTD by eliminating all site-specific parameters from the SPINN-TD empirical equation (Equation 8), as outlined in Equation 9:
${\hat{y}}_{st}=p_{1}\left(y_{1_{max}}\right)^{-\alpha}\left(q_{max}\right)^{\beta}\left(1-e^{-}\frac{t}{T_{L}}\right)$ (9)
@@
This approach enables the SPINN model to get trained with a much larger database and learn the features of a wider range of scour events from a cluster of bridges showing similar flow and scour trends. Such transferable model is envisaged to be able to collate and transfer the learning among several bridges. Transferable models are particularly relevant for new bridges with limited data available. The calibrated general time-dependent (GTD) equation may be viable to provide reasonable estimates of maximum scour depth in a new bridge within the same region/cluster. We evaluated this hypothesis in the following sections.
2.3.4 Base DL Algorithm: 1-Long Short-Term Memory Networks (LSTM)